# Hitting probabilities for fast stochastic search

### v1

@@
SL: Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (samantha.linn@utah.edu). SL was supported by the National Science Foundation (Grant No.  2139322)
@@ 1. Introduction
In the fast search limit, only an approximation of the short-time behavior of the unconditional search process is needed to compute the conditional hitting probabilities. To be precise, consider an unconditional search process and let $\{X(t)\}_{t\geq 0}$ denote the path of the stochastic searcher in a domain $\mathcal{D}\in\mathbb{R}^{d}$ with $m\geq 2$ targets denoted by $V_{0}$, $V_{1}$, $\dots$, $V_{m-1}\in\mathcal{D}$. Then the first hitting time to a target is
$\tau:=\text{inf}\{t>0:X(t)\in\cup_{k=0}^{m-1}V_{k}\in\mathcal{D}\}.$
Let $\kappa\in\{0,1,\dots,m-1\}$ denote the index of the target hit by the searcher. Then the hitting probability of target $V_{k}$ is
$\mathbb{P}(\kappa=k)\quad\text{for }k\in\{0,1,\dots,m-1\}.$
@@
Figure 1. Hitting probabilities for diffusion with unit diffusivity, initial position $x_{0}\in(0,\ell/2)$. Solid colored curves illustrate case of exponentially distributed resetting times $\sigma>0$ with rate $r>0$; dashed black curve denotes no conditioning.
As a motivating example, consider a driftless diffusive search process with diffusivity $D>0$ in one dimension. Without loss of generality, assume the searcher starts at $x_{0}\in(0,\ell/2)$ where $V_{0}=(-\infty,0]$ and $V_{1}=[\ell,\infty)$ are targets. Let $L_{0}>0$ and $L_{1}>0$ be the distances to $V_{0}$ and $V_{1}$ from $x_{0}$, respectively. In the case that there is no condition on the search time, one can show [45] that the probability of the searcher reaching $V_{1}$ before $V_{0}$ is
@@
Now suppose we impose resetting on the searcher, which we denote by the random variable $\sigma>0$. If $\sigma>0$ is exponentially distributed with rate $r>0$, we show that the hitting probability to $V_{1}$ decays to zero like
@@
(2) $\mathbb{P}(\kappa=k|\tau<\sigma)\quad\text{for }k\in\{0,1,\dots,m-1\}.$
@@ 2.1. Probabilistic setup and integral representation
Let $\tau>0$ be a strictly positive random variable and let $\kappa$ be a random variable taking values from the set $\{0,1,\dots,m-1\}$. In the applications of this text, $\tau>0$ denotes the hitting time of an unconditional search process to a target and $\kappa$ indexes which of the $m\geq 2$ targets is hit. For each index $k\in\{0,1,\dots,m-1\}$, let $\tau^{(k)}$ be the hitting time to target $V_{k}$,
@@
$F_{\tau}(t):=\mathbb{P}(\tau\leq t)=\sum_{k=0}^{m-1}F_{\tau}^{(k)}(t),\quad t\in\mathbb{R}.$
@@
where we call the parameter $r>0$ the resetting rate and $Y>0$ is a strictly positive random variable with unit mean and a finite moment generating function in a neighborhood of the origin. That is,
@@
The second condition in (4) implies that $S_{Y}(y):=\mathbb{P}(Y>y)$ decays at least exponentially fast,
@@ 2.2. Exact hitting probability asymptotics
Assume $C,\delta>0$ and $b\in\mathbb{R}$. Then
@@
Under the assumptions of section 2.1, assume further that $\sigma>0$ is gamma distributed with shape $\eta>0$ and rate $r>0$ and that for some $k\in\{1,\dots,m-1\}$,
@@
where $A>0$, $B>0$, $C>0$, $C_{k}>0$, and $m,\,n\in\mathbb{R}$. Then
@@
We remark that in both Theorem 2 and Corollary 3 the dominant behavior of the asymptotic hitting probabilities depends only on $C$ and $C_{k}$. We show in section 3 that this dependence is in fact only on the geodesic lengths to the nearest target and to target $k$ for diffusive search. Also observe that as $r$ increases in Theorem 2 and Corollary 3, the asymptotic hitting probabilities behave progressively more like decaying exponential functions. The same qualitative behavior is true in the following theorem, which is relevant to typical scenarios of diffusive search under uniformly distributed resetting.
@@
Under the assumptions of section 2.1, assume further that $\sigma>0$ is uniformly distributed on $[0,2/r]$ and for some $k\in\{1,\dots,m-1\}$,
@@
where $A>0$, $B>0$, $C>0$, $C_{k}>0$, and $m,\,n\in\mathbb{R}$. Then
@@
Under the assumptions of section 2.1, assume further that $\sigma>0$ is gamma distributed with shape $\eta>0$ and rate $r>0$ and that for some $k\in\{1,\dots,m-1\}$,
@@
The following corollary, while immediately clear from Theorem 5, applies to superdiffusive search [46] and fundamentally differs from instances of diffusive search. In particular, these probabilities are between zero and one for every target. In other words, it is not the case that the nearest target is found in the large $r$ limit.
@@
Under the assumptions of section 2.1, assume further that the short-time behavior of $F_{\tau}$ and $F_{\tau}^{(k)}$ are known,
@@
$F_{\tau}^{(k)}(t)$ $\sim\tilde{F}_{\tau}^{(k)}(t)\quad\text{as }t\to 0^{+},$
and that $\sigma>0$ is deterministic and equal to $\sigma:=1/r$. Then
$\mathbb{P}(\kappa=k|\tau<\sigma)\sim\frac{\tilde{F}_{\tau}^{(k)}(1/r)}{\tilde{F}_{\tau}(1/r)}\quad\text{as }r\to\infty.$
@@ 3.1. Pure diffusion in one dimension
Figure 2. Conditional hitting probabilities for diffusion with diffusivity $D=1$, drift $\mu\in\mathbb{R}$, initial position $x_{0}\in(0,\ell)$, and exponentially distributed $\sigma>0$ with rate $r>0$. Curves illustrate analytical asymptotic estimates; markers denote numerical solutions. See subsections 3.1 and 3.2 for details.
@@
$\mathbb{P}(\kappa=1|\tau<\sigma)\sim e^{-\sqrt{\frac{(L_{1}-L_{0})r}{D}}}=e^{-\sqrt{\frac{(\ell-2x_{0})^{2}r}{D}}}\quad\text{as }r\to\infty.$
@@ 3.3. Diffusion with partially absorbing targets
$\tau_{\text{partial}}:=\inf\{t>0:\lambda_{k}>\xi_{k}/\gamma_{k}\text{ for some }k\in\{0,\dots,m-1\}\}$
where $\{\xi_{k}\}_{k=0}^{m-1}$ are unit rate exponential random variables, $\{\gamma_{k}\}_{k=0}^{m-1}$ are non-negative trapping rates, and $\lambda_{k}$ is the local time of $X(t)$ on target $V_{k}$ [48].
@@
The asymptotic conditional hitting probabilities for pure diffusion with partially absorbing targets in the interval with numerous distributions for $\sigma$ are catalogued in Table 1. Figure 3 illustrates these probabilities and compares them to numerical solutions.
@@ 3.4. Diffusion between concentric spheres
Figure 3. Left: Conditional hitting probabilities for diffusion with diffusivity $D>0$ between concentric spheres of radii $R_{0}$ and $R_{1}$ and exponentially distributed $\sigma>0$ with rate $r>0$. Above, $y_{0}:=(x_{0}-R_{0})/(R_{1}-R_{0})$. Curves illustrate analytical solutions; markers denote numerical solutions. See subsection 3.4 for details. Right: Conditional hitting probabilities for an RTP with velocity $v>0$, tumble rate $\lambda>0$, and exponentially distributed $\sigma>0$ with rate $r>0$. Curves illustrate analytical solutions; markers denote simulations. See subsection 3.7 for details.
@@ 3.5. Random walk on a network
Now consider the spatially discrete stochastic search process of a random walk on a discrete network. To be precise, let $\{X(t)\}_{t\geq 0}$ be a continuous-time Markov chain on a finite or countably infinite state space, $I$. We assume the jump rates are bounded to preclude the possibility of infinitely many jumps occurring in a short time on a countably infinite network. Denote by $i_{0}\in I$ the initial position of $X(t)$ and denote target nodes (or sets of nodes that constitute targets) by $V_{k}\in I$ where $i_{0}\not\in\cup_{k=0}^{m-1}V_{k}$.
@@ 3.6. Superdiffusive Lévy flight in one dimension
Suppose now the stochastic search process is a superdiffusive Lévy flight starting from $x_{0}\in(0,\ell/2)$ where $x=0$ and $x=\ell$ denote targets and we are interested in the first passage time of the search over each target. Equivalently, we may denote the targets by $V_{0}=(-\infty,0]$ and $V_{1}=[\ell,\infty)$ and study the first hitting time to either target.
@@
$\partial_{t}p=-K(-\Delta)^{\alpha/2}p,\quad t>0,$
where $K>0$ is the generalized diffusivity and $(-\Delta)^{\alpha/2}$ is the fractional Laplacian with $\alpha\in(0,2)$ [50]. Assuming $x_{0}\notin V_{0}\cup V_{1}$, one can show that $F_{\tau}$ and $F_{\tau}^{(1)}$ satisfy (12) and (13) with
$A=\frac{\Gamma(\alpha)K\sin(\alpha\pi/2)}{\pi(\ell-x_{0})^{\alpha}}+\frac{\Gamma(\alpha)K\sin(\alpha\pi/2)}{\pi x_{0}^{\alpha}},\quad B=\frac{\Gamma(\alpha)K\sin(\alpha\pi/2)}{\pi(\ell-x_{0})^{\alpha}}$

### v2

@@
SL (corresponding author): Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (slinn.math@gmail.com). SL was supported by the National Science Foundation (Grant No.  2139322)
@@ 1. Introduction
In the fast search limit, only an approximation of the short-time behavior of the unconditional search process is needed to compute the conditional hitting probabilities. To be precise, consider an unconditional search process and let $\{X(t)\}_{t\geq 0}$ denote the path of the stochastic searcher in a domain $\mathcal{D}\in\mathbb{R}^{d}$ with $K\geq 2$ targets denoted by $V_{0}$, $V_{1}$, $\dots$, $V_{K-1}\in\mathcal{D}$. Then the first hitting time to a target is
$\tau:=\text{inf}\{t>0:X(t)\in\cup_{k=0}^{K-1}V_{k}\in\mathcal{D}\}.$
Let $\kappa\in\{0,1,\dots,K-1\}$ denote the index of the target hit by the searcher. Then the hitting probability of target $V_{k}$ is
$\mathbb{P}(\kappa=k)\quad\text{for }k\in\{0,1,\dots,K-1\}.$
@@
Figure 1. Hitting probabilities for diffusion with unit diffusivity, initial position $x_{0}\in(0,\ell)$. Solid colored curves illustrate case of exponentially distributed resetting times $\sigma>0$ with rate $r>0$; dashed black curve denotes no conditioning.
As a motivating example, consider a driftless diffusive search process with diffusivity $D>0$ in one dimension. Assume the searcher starts at $x_{0}\in(0,\ell/2)$ where $V_{0}=(-\infty,0]$ and $V_{1}=[\ell,\infty)$ are targets. Let $L_{0}>0$ and $L_{1}>0$ be the distances to $V_{0}$ and $V_{1}$ from $x_{0}$, respectively. In the case that there is no condition on the search time, one can show [45] that the probability of the searcher reaching $V_{1}$ before $V_{0}$ is
@@
Now suppose the searcher stochastically resets to its initial position $x_{0}$ at random independent and identically-distributed times governed by $\sigma>0$. If $\sigma>0$ is exponentially distributed with rate $r>0$, we show in section 3.1 that the hitting probability to $V_{1}$ decays to zero like
@@
(2) $\mathbb{P}(\kappa=k|\tau<\sigma)\quad\text{for }k\in\{0,1,\dots,K-1\}.$
@@ 2.1. Probabilistic setup and integral representation
Let $\tau>0$ be a strictly positive random variable and let $\kappa$ be a random variable taking values from the set $\{0,1,\dots,K-1\}$. In the applications of this text, $\tau>0$ denotes the hitting time of an unconditional search process to a target and $\kappa$ indexes which of the $K\geq 2$ targets is hit. For each index $k\in\{0,1,\dots,K-1\}$, let $\tau^{(k)}$ be the hitting time to target $V_{k}$,
@@
$F_{\tau}(t):=\mathbb{P}(\tau\leq t)=\sum_{k=0}^{K-1}F_{\tau}^{(k)}(t),\quad t\in\mathbb{R}.$
@@
where we call the parameter $r>0$ the resetting rate and $Y>0$ is a strictly positive random variable with unit mean and a finite moment generating function in a neighborhood of the origin. That is, there exists $\eta>0$ such that
@@
Applying the second condition in (4) and Chebyshev’s inequality (see, for example, Theorem 1.6.4 in [46]), we obtain that $S_{Y}(y):=\mathbb{P}(Y>y)$ decays at least exponentially fast,
@@ 2.2. Exact hitting probability asymptotics
Assume $C>0$, $\delta>0$ and $b\in\mathbb{R}$. Then
@@
Under the assumptions of section 2.1, assume further that $\sigma>0$ is gamma distributed with shape $\eta>0$ and rate $r>0$ and that for some $k\in\{1,\dots,K-1\}$,
@@
where $A>0$, $B>0$, $C_{k}>C>0$, and $m,\,n\in\mathbb{R}$. Then
@@
We remark on the exclusion of $k=0$ from Theorem 2 above: Information about the nearest target determines the constants in (9). Hence if $k=0$ the expressions in (9) and (10) are identical and the asymptotic hitting probability is trivial. For this reason, $k=0$ is also omitted from the remaining results that follow.
Moreover, we note that in both Theorem 2 and Corollary 3 the dominant behavior of the asymptotic hitting probabilities depends only on $C$ and $C_{k}$. We show in section 3 that this dependence is in fact only on the geodesic lengths to the nearest target and to target $k$ for diffusive search. Also observe that as $r$ increases in Theorem 2 and Corollary 3, the asymptotic hitting probabilities behave progressively more like decaying exponential functions. The same qualitative behavior is true in the following theorem, which is relevant to typical scenarios of diffusive search under uniformly distributed resetting.
@@
Under the assumptions of section 2.1, assume further that $\sigma>0$ is uniformly distributed on $[0,2/r]$ and for some $k\in\{1,\dots,K-1\}$,
@@
where $A>0$, $B>0$, $C_{k}>C>0$, and $m,\,n\in\mathbb{R}$. Then
@@
Under the assumptions of section 2.1, assume further that $\sigma>0$ is gamma distributed with shape $\eta>0$ and rate $r>0$ and that for some $k\in\{1,\dots,K-1\}$,
@@
The following corollary, while immediately clear from Theorem 5, applies to superdiffusive search [47] and fundamentally differs from instances of diffusive search. In particular, these probabilities are neither zero nor one for every target in the asymptotic limit, hence, the nearest target may not be found.
@@
Under the assumptions of section 2.1, assume further that $\sigma>0$ is deterministic and equal to $\sigma:=1/r$. Then
$\mathbb{P}(\kappa=k|\tau<\sigma)=\frac{F_{\tau}^{(k)}(1/r)}{F_{\tau}(1/r)}\sim\frac{\tilde{F}_{\tau}^{(k)}(1/r)}{\tilde{F}_{\tau}(1/r)}\quad\text{as }r\to\infty,$
where $\tilde{F}_{\tau}$ and $\tilde{F}_{\tau}^{(k)}$ denote the short-time behavior of $F_{\tau}$ and $F_{\tau}^{(k)}$, respectively,
@@
$F_{\tau}^{(k)}(t)$ $\sim\tilde{F}_{\tau}^{(k)}(t)\quad\text{as }t\to 0^{+}.$
@@ 3.1. Pure diffusion in one dimension
Figure 2. Conditional hitting probabilities for diffusion with diffusivity $D=1$, drift $\mu\in\mathbb{R}$, initial position $x_{0}\in(0,\ell/2)$, and exponentially distributed $\sigma>0$ with rate $r>0$. Curves illustrate analytical asymptotic estimates; markers denote numerical solutions. See subsections 3.1 and 3.2 for details.
@@
$\mathbb{P}(\kappa=1|\tau<\sigma)\sim e^{-\sqrt{\frac{(L_{1}-L_{0})^{2}r}{D}}}=e^{-\sqrt{\frac{(\ell-2x_{0})^{2}r}{D}}}\quad\text{as }r\to\infty.$
@@ 3.3. Diffusion with partially absorbing targets
$\tau_{\text{partial}}:=\inf\{t>0:\lambda_{k}>\xi_{k}/\gamma_{k}\text{ for some }k\in\{0,\dots,K-1\}\}$
where $\{\xi_{k}\}_{k=0}^{K-1}$ are unit rate exponential random variables, $\{\gamma_{k}\}_{k=0}^{K-1}$ are non-negative trapping rates, and $\lambda_{k}$ is the local time of $X(t)$ on target $V_{k}$ [49].
@@
Figure 3. Left: Conditional hitting probabilities for diffusion with diffusivity $D>0$ between concentric spheres of radii $R_{0}$ and $R_{1}$ and exponentially distributed $\sigma>0$ with rate $r>0$. Above, $y_{0}:=(x_{0}-R_{0})/(R_{1}-R_{0})$. Curves illustrate analytical asymptotic estimates; markers denote numerical solutions. See subsection 3.4 for details. Right: Conditional hitting probabilities for an RTP with velocity $v>0$, tumble rate $\lambda>0$, and exponentially distributed $\sigma>0$ with rate $r>0$. Curves illustrate analytical asymptotic estimates; markers denote simulations. See subsection 3.7 for details.
@@
The asymptotic conditional hitting probabilities for pure diffusion with partially absorbing targets in the interval with numerous distributions for $\sigma$ are catalogued in Table 1.
@@ 3.5. Random walk on a network
Now consider the spatially discrete stochastic search process of a random walk on a discrete network. To be precise, let $\{X(t)\}_{t\geq 0}$ be a continuous-time Markov chain on a finite or countably infinite state space, $I$. We assume the jump rates are bounded to preclude the possibility of infinitely many jumps occurring in a short time on a countably infinite network. Denote by $i_{0}\in I$ the initial position of $X(t)$ and denote target nodes (or sets of nodes that constitute targets) by $V_{k}\in I$ where $i_{0}\not\in\cup_{k=0}^{K-1}V_{k}$.
@@
Figure 4. Left: Path of an unconditional $\alpha=1$ Lévy flight search. Markers denote resetting. Right: Conditional hitting probabilities for Lévy flights with $\alpha\in(0,2)$ and exponentially distributed $\sigma>0$ with rate $r>0$. Lines illustrate analytical asymptotic estimates; markers denote simulations. See subsection 3.6 for details.
@@ 3.6. Superdiffusive Lévy flight in one dimension
Suppose now the stochastic search process is a superdiffusive Lévy flight starting from $x_{0}\in(0,\ell/2)$ where $V_{0}=(-\infty,0]$ and $V_{1}=[\ell,\infty)$ are targets.
@@
$\partial_{t}p=-\bar{D}(-\Delta)^{\alpha/2}p,\quad t>0,$
where $\bar{D}>0$ is the generalized diffusivity and $(-\Delta)^{\alpha/2}$ is the fractional Laplacian with $\alpha\in(0,2)$ [51]. Assuming $x_{0}\notin V_{0}\cup V_{1}$, one can show that $F_{\tau}$ and $F_{\tau}^{(1)}$ satisfy (12) and (13) with
$A=\frac{\Gamma(\alpha)\bar{D}\sin(\alpha\pi/2)}{\pi(\ell-x_{0})^{\alpha}}+\frac{\Gamma(\alpha)\bar{D}\sin(\alpha\pi/2)}{\pi x_{0}^{\alpha}},\quad B=\frac{\Gamma(\alpha)\bar{D}\sin(\alpha\pi/2)}{\pi(\ell-x_{0})^{\alpha}}$
@@ 4. Application to a decision-making model
While such hasty decisions may not be made frequently by a single agent, first deciders among groups of agents are reliably quick and similarly biased [57]. The influence of these early decisions on agents still deliberating can be substantial; recent work suggests that having many fast deciders can sway the group to the same hasty decision, but having only a few fast deciders can caution remaining agents against biased decision-making [34].
@@ 6.3.3. Run-and-tumble in one dimension
6.3.4. Superdiffusive Lévy flight in one dimension
In Figure 4 we illustrate conditional hitting probabilities corresponding to a superdiffusive Lévy flight in one dimension as detailed in subsection 3.6. We now describe the stochastic simulation algorithm used herein. Given a discrete time step $\Delta t>0$, we generate a statistically exact path of the $(\alpha/2)$-subordinator $U=\{U(t)\}_{t\geq 0}$ on the discrete time grid $\{t_{n}\}_{n\in\mathbb{N}}$ with $t_{n}=n\Delta t$ via
$U(t_{n+1})=U(t_{n})+(\Delta t)^{2/\alpha}\Theta_{n},\quad n\geq 0,$
where $U(t_{0})=U(0)=0$ and $\{\Theta_{n}\}_{n\in\mathbb{N}}$ is an independent and identically-distributed sequence of realizations
$\Theta=\frac{\sin((\alpha/2)(V+\pi/2))}{(\cos V)^{2/\alpha}}\left(\frac{\cos(V-(\alpha/2)(V+\pi/2))}{E}\right)^{(2-\alpha)/\alpha},$
where $V$ is uniformly distributed on $(-\pi/2,\pi/2)$ and $E$ is an independent unit mean exponential random variable [68]. From this we can generate a statistically exact path of Brownian motion $\{B(s)\}_{s\geq 0}$ on the discrete time grid $\{U(t_{n})\}_{n\in\mathbb{N}}$ via
$B(U(t_{n+1}))=B(U(t_{n}))+\sqrt{2(D_{s}\Delta t)^{2/\alpha}\Theta_{n}}\xi_{n},\quad n\geq 0,$
where $\{\xi_{n}\}_{n\in\mathbb{N}}$ is an independent and identically-distributed sequence of standard Gaussian random variables. Finally one obtains a statistically exact path of a Lévy flight with generalized diffusivity $D_{s}>0$ via the following random time change of the Brownian motion $B$,
$X(t):=D_{s}^{1/\alpha}B(U(t)),\quad t\geq 0.$
To determine the conditional hitting probabilities, we generate an exponentially distributed resetting time and simulate a Lévy flight path until the reset time. If the path reaches a target prior to the resetting time, the process ends and we note which target was found. Otherwise we reset the search to its initial position and start the process anew. Repeating this process for a given resetting rate allows us to compute the corresponding conditional hitting probability. Each data point in Figure 4 is computed from 12000 independent trials with $D_{s}=1$ and $\Delta t=10^{-4}$.
Acknowledgements
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2139322. SDL was supported by the National Science Foundation (Grant Nos. DMS-2325258 and DMS-1944574). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.